# Treewidth 2 in the Planar Graph Product Structure Theorem

### v1

@@ 1 Introduction
As illustrated in Figure 1, the strong product $A\mathbin{\boxtimes}B$ of graphs $A$ and $B$ has vertex-set $V(A)\times V(B)$, where distinct vertices $(v,x),(w,y)$ are adjacent if:
@@
Theorem 1 provides a powerful tool for studying questions about planar graphs, by reducing to graphs of bounded treewidth. Indeed, this result has been the key for resolving several open problems regarding queue layouts [8], nonrepetitive colourings [7], centred colourings [5], adjacency labelling schemes [11, 2, 10, 6], twin-width [3, 15], infinite graphs [13], and comparable box dimension [9]. In several of these applications, because the dependence on $\operatorname{tw}(H)$ is often exponential, the best bounds are obtained by applying the 3-term product in Theorem 1(c).
@@ 3 Proof of Theorem 2
If every planar graph had a vertex-partition into two induced forests, then the Four-Colour Theorem would follow. Chartrand and Kronk [4] constructed planar graphs that have no vertex-partition into two induced forests. On the other hand, Thomassen [19, Theorem 4.1] showed the following analogous result where the forest requirement is relaxed [^1] Thomassen [19, Theorem 4.1] proved Lemma 5 for planar triangulations, which implies the result for general planar graphs, since any subgraph of a triangle-forest is a triangle-forest. Lemma 5 was rediscovered by Knauer et al. [16].. A triangle-forest is a graph in which every cycle is a triangle.
@@
We may assume that $G$ has an edge $ab$ such that $G^{\prime}:=(G-a)-b$ is planar. By Corollaries 7 and 4, $G^{\prime}$ has a matching $M^{\prime}$ such that $(G^{\prime}/M^{\prime})^{+}\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2$ for some apex-forests $H_{\_}1$ and $H_{\_}2$. Thus, for the matching $M:=M^{\prime}\cup\{ab\}$ we have $G/M\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}(G^{\prime}/M^{\prime})^{+}\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2$. Hence $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2\boxtimes K_{\_}2$. Here we use the fact that if $M$ is a matching in a graph $G$, then $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}(G/M)\boxtimes K_{\_}2$. ∎
@@ 4 Proof of Theorem 3
()($\Leftarrow$)( ⇐ ) Assume $G$ has partitions $\mathcal{P}_{\_}1$ and $\mathcal{P}_{\_}2$ such that $G/\mathcal{P}_{\_}i\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}i$ for each $i\in\{1,2\}$, and $|A_{\_}1\cap A_{\_}2|\leqslant c$ for each $A_{\_}1\in\mathcal{P}_{\_}1$ and $A_{\_}2\in\mathcal{P}_{\_}2$. Let $\phi_{\_}i$ be an isomorphism from $G/\mathcal{P}_{\_}i$ to a subgraph of $H_{\_}1$. For each $A_{\_}1\in\mathcal{P}_{\_}1$ and $A_{\_}2\in\mathcal{P}_{\_}2$, enumerate the at most $c$ vertices in $A_{\_}1\cap A_{\_}2$, and map the $i$-th such vertex to $(\phi_{\_}1(A_{\_}1),\phi_{\_}2(A_{\_}2),i)$. This defines an isomorphism from $G$ to a subgraph of $H_{\_}1\mathbin{\boxtimes}H_{\_}2\mathbin{\boxtimes}K_{\_}c$, as desired. ∎
@@
For any $c\in\mathbb{N}$, let $F$ be a double-fan on at least $18c^{2}+2c+$ 1 vertices. Let $\mathcal{P},\mathcal{Q}$ be partitions of $F$ such that $\mathcal{P}$ is a tree-partition and $|P\cap Q|\leqslant c$ for all $P\in\mathcal{P}$ and $Q\in\mathcal{Q}$. Let $v_{\_}1,v_{\_}2$ be the centres of $F$. If $P_{\_}1\neq P_{\_}2$ and $Q_{\_}1\neq Q_{\_}2$, then there exists a 4-clique $\{v_{\_}1,v_{\_}2,v_{\_}3,v_{\_}4\}$ with $Q_{\_}1,Q_{\_}2,Q_{\_}3,Q_{\_}4$ pairwise distinct.
@@
Since $|V(F)|=t+1\geqslant c^{2}+c+1$, by Lemma 10 applied to $F$ and the induced partitions of $F$, there exists $v_{\_}2\in B$ such that $Q_{\_}1\neq Q_{\_}2$. Let $C$ be the $t$-vertex path $J_{\_}{v_{\_}1v_{\_}2}-\{v_{\_}1,v_{\_}2\}$.
Since $|V(J_{\_}{v_{\_}1v_{\_}2})|=t+2=8c^{2}+2c+1$, if $P_{\_}1\neq P_{\_}2$ then by Lemma 11, there exists a 4-clique $\{v_{\_}1,v_{\_}2,v_{\_}3,v_{\_}4\}$ in $J_{\_}{v_{\_}1v_{\_}2}$ with $Q_{\_}1,Q_{\_}2,Q_{\_}3,Q_{\_}4$ pairwise distinct, as desired. So we may assume that $P_{\_}1=P_{\_}2$.

### v2

@@ 1 Introduction
As illustrated in Figure 1, the strong product $A\mathbin{\boxtimes}B$ of graphs $A$ and $B$ has vertex-set
Figure 1: Strong product of paths.
$V(A)\times V(B)$, where distinct vertices $(v,x),(w,y)$ are adjacent if:
@@
Theorem 1 provides a powerful tool for studying questions about planar graphs, by reducing to graphs of bounded treewidth. Indeed, this result has been the key for resolving several open problems regarding queue layouts [8], nonrepetitive colourings [7], centred colourings [5], adjacency labelling schemes [11, 2, 10, 6], twin-width [3, 17, 15], infinite graphs [13], and comparable box dimension [9]. In several of these applications, because the dependence on $\operatorname{tw}(H)$ is often exponential, the best bounds are obtained by applying the 3-term product in Theorem 1(c).
@@
We conclude this introduction by mentioning an open problem: Does Theorem 2 hold with $H_{\_}1\boxtimes H_{\_}2\boxtimes K_{\_}2$ replaced by $H_{\_}1\boxtimes H_{\_}2$?
@@ 3 Proof of Theorem 2
If every planar graph had a vertex-partition into two induced forests, then the Four-Colour Theorem would follow. Chartrand and Kronk [4] constructed planar graphs that have no vertex-partition into two induced forests. On the other hand, Thomassen [20, Theorem 4.1] showed the following analogous result where the forest requirement is relaxed [^1] Thomassen [20, Theorem 4.1] proved Lemma 5 for planar triangulations, which implies the result for general planar graphs, since any subgraph of a triangle-forest is a triangle-forest. Lemma 5 was rediscovered by Knauer et al. [16].. A triangle-forest is a graph in which every cycle is a triangle.
@@
We may assume that $G$ has an edge $ab$ such that $G^{\prime}:=G-a-b$ is planar. By Corollaries 7 and 4, $G^{\prime}$ has a matching $M^{\prime}$ such that $(G^{\prime}/M^{\prime})^{+}\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2$ for some apex-forests $H_{\_}1$ and $H_{\_}2$. Thus, for the matching $M:=M^{\prime}\cup\{ab\}$ we have $G/M\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}(G^{\prime}/M^{\prime})^{+}\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2$. Hence $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2\boxtimes K_{\_}2$. Here we use the fact that if $M$ is a matching in a graph $G$, then $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}(G/M)\boxtimes K_{\_}2$. ∎
@@ 4 Proof of Theorem 3
()($\Leftarrow$)( ⇐ ) Assume $G$ has partitions $\mathcal{P}_{\_}1$ and $\mathcal{P}_{\_}2$ such that $G/\mathcal{P}_{\_}i\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}i$ for each $i\in\{1,2\}$, and $|A_{\_}1\cap A_{\_}2|\leqslant c$ for each $A_{\_}1\in\mathcal{P}_{\_}1$ and $A_{\_}2\in\mathcal{P}_{\_}2$. Let $\phi_{\_}i$ be an isomorphism from $G/\mathcal{P}_{\_}i$ to a subgraph of $H_{\_}i$. For each $A_{\_}1\in\mathcal{P}_{\_}1$ and $A_{\_}2\in\mathcal{P}_{\_}2$, enumerate the at most $c$ vertices in $A_{\_}1\cap A_{\_}2$, and map the $i$-th such vertex to $(\phi_{\_}1(A_{\_}1),\phi_{\_}2(A_{\_}2),i)$. This defines an isomorphism from $G$ to a subgraph of $H_{\_}1\mathbin{\boxtimes}H_{\_}2\mathbin{\boxtimes}K_{\_}c$, as desired. ∎
@@
For any $c\in\mathbb{N}$, let $F$ be a double-fan on at least $18c^{2}+2c+$ 1 vertices. Let $\mathcal{P},\mathcal{Q}$ be partitions of $F$ such that $\mathcal{P}$ is a tree-partition and $|P\cap Q|\leqslant c$ for all $P\in\mathcal{P}$ and $Q\in\mathcal{Q}$. Let $v_{\_}1,v_{\_}2$ be the centres of $F$. If $P_{\_}1\neq P_{\_}2$ and $Q_{\_}1\neq Q_{\_}2$, then there exist vertices $v_{\_}3,v_{\_}4$ such that $\{v_{\_}1,v_{\_}2,v_{\_}3,v_{\_}4\}$ is a $4$-clique and $Q_{\_}1,Q_{\_}2,Q_{\_}3,Q_{\_}4$ are pairwise distinct.
@@
Since $|V(F)|=t+1\geqslant c^{2}+c+1$, by Lemma 10 applied to $F$ and the induced partitions of $F$, there exists $v_{\_}2\in B$ such that $Q_{\_}1\neq Q_{\_}2$. Let $C$ be the $t$-vertex path $J_{\_}{v_{\_}1v_{\_}2}-v_{\_}1-v_{\_}2$.
Since $|V(J_{\_}{v_{\_}1v_{\_}2})|=t+2=8c^{2}+2c+1$, if $P_{\_}1\neq P_{\_}2$ then by Lemma 11, there exist vertices $v_{\_}3,v_{\_}4$ such that $\{v_{\_}1,v_{\_}2,v_{\_}3,v_{\_}4\}$ is a $4$-clique $J_{\_}{v_{\_}1v_{\_}2}$ with $Q_{\_}1,Q_{\_}2,Q_{\_}3,Q_{\_}4$ pairwise distinct, as desired. So we may assume that $P_{\_}1=P_{\_}2$.